OpenStudy (anonymous):

ook now this one find the radius of the circle defined by the equation x^+y^2-10x+4y+4=0

OpenStudy (anonymous):

i found it i got a different one

OpenStudy (cherylim23):

Complete the square to first get the equation of the circle. x^2+y^2-10x+4y+4=0 (You forgot a power 2 somewhere there) (x-5)^2 - 5^2 + (y+2)^2 -2^2 + 4 = 0 (x-5)^2 + (y+2)^2 = 25-4 + 4 (x-5)^2 + (y+2)^2 = 25 Centre is thus (5,-2) and the radius, which you have already found, should be sqrt (25) = 5 units

OpenStudy (anonymous):

\[x ^{2}-10x+y ^{2}+4y+4=0\] \[x ^{2}-10x+25+y ^{2}+4y+4=25\] \[(x-5)^{2}+(y+2)^{2}=25\]

OpenStudy (anonymous):

ook now this one find the radius of the circle defined by the equation 3x^2+3y^2+42x+42y+186=0

OpenStudy (anonymous):

divide 3 on both sides firstly

OpenStudy (anonymous):

um hw

OpenStudy (anonymous):

\[(3x ^{2}+3y^{2}+42x+42y+186)/3=0/3\]

OpenStudy (anonymous):

\[x^{2}+y^{2}+14x+14y+62=0\]

OpenStudy (anonymous):

ok now tell me what to do

OpenStudy (anonymous):

\[(x+a)^2+(x+b)^2=c\] a,b,c\[a,b,c \in R\]

OpenStudy (anonymous):

um whats a b c

OpenStudy (anonymous):

for (x−5)^2+(y+2)^2=25 a=-5,b=2,c=25

OpenStudy (anonymous):

um-337? idk this is hard

OpenStudy (anonymous):

what is your a ,b and c values?

OpenStudy (anonymous):

idk -5 2 25

OpenStudy (anonymous):

\[r=\sqrt{c}\]

OpenStudy (anonymous):

so is it 5 sooooooo is it 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you solve the 3x^2+3y^2+42x+42y+186=0 on you own now ?Plz have a try